• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 14:53:02 on October 7,2002

    Entry number 87564

    This entry is a followup to: 87555

    keyword=charts etc.

    Chris Keith fixed the problems with chart display: now one can change scale.
    Vacuum reading is fine. Beam current is using the Hall A upstream BCM now.

    Also I changed the alarm setpoints for LD2 loop and saved to bootfile.

    Fan frequency is 24 Hz. And should be set to 60 Hz before beam arrives
    (next shift). HP heat load is about 100 W now, need to be adjusted before
    beam arrives.